OpenStudy (anonymous):

h(x)=5x^4-3x^2-2 h(-x)=? I got 2x-2 for the answer, I don't get what I did wrong

OpenStudy (freckles):

You are suppose to have the same amount of terms in both h(x) and h(-x)

OpenStudy (freckles):

replace all the x's in h(x) with -x

OpenStudy (freckles):

now this can be cleaned up a little

OpenStudy (freckles):

\[h(-x)=5(-x)^4-3(-x)^2-2*\]

OpenStudy (freckles):

try to simplify (-x)^4

OpenStudy (anonymous):

=x

OpenStudy (freckles):

well not exactly...

OpenStudy (freckles):

\[(-x)^4=(-1 \cdot x)^4=(-1)^4 \cdot x^4=?\]

OpenStudy (anonymous):

1x?

OpenStudy (anonymous):

x^4

OpenStudy (freckles):

yes x^4

OpenStudy (freckles):

and what does(-x)^2=?

OpenStudy (anonymous):

x^2

OpenStudy (freckles):

\[h(-x)=5(-x)^4-3(-x)^2-2 \\ h(-x)=5(-1 \cdot x)^4-3(-1 \cdot x)^2-2 \\ h(-x)=5(-1)^4(x)^4-3(-1)^2(x)^2-2\\h(-x)=5(1)x^4-3(1)x^2-2\\h(-x)=5x^4-3x^2-2\]

OpenStudy (freckles):

So in this case h(-x) actually equals h(x)

OpenStudy (freckles):

Which implies a certain kinda symmetry

OpenStudy (freckles):

But the question was just what is h(-x)

OpenStudy (anonymous):

hmm thats confusing, but i see what you did. thank you!

OpenStudy (freckles):

what part confuses you?

OpenStudy (freckles):

Could you find h(-x) if h(x)=x^3-5x^2+x-1

OpenStudy (anonymous):

yea, i should have looked at it the same way

OpenStudy (freckles):

yes except here you actually have odd powers so ... if you have \[(-x)^{even power}=x^{evenpower} \\ (-x)^{oddpower}=-x^{oddpower}\]

OpenStudy (freckles):

That will be a good thing to know for this type of thing

OpenStudy (freckles):

So what I'm saying is example of that first thing I said \[(-x)^{20}=x^{20}\] example of the second thing I said \[(-x)^{207}=-x^{207}\]

OpenStudy (freckles):

Notice the 20 was even and the 207 was odd

OpenStudy (anonymous):

oh okay, i see

OpenStudy (freckles):

Let me know when you have found h(-x) for this new h(x) I made for you